Evaluate the exponential equation for three values of x greater than -4, three values of x smaller than -4, and at x = -4.
y = 3(x + 4)

Answer by checkley77(12844)

y = 3(x + 4)
Let x=-3,5,7
y=3(-3+4)=3*1=3
y=3(5+4)=3*9=27
y=3(7+4)=3*11=33
---------------
Let x=-5,-7,-9
y=3(-5+4)=3*-1=-3
y=3(-7+4)=3*-3=-9
y=3(-9+4)=3*-5=-15
--------------
Let x=-4
y=3(-4+4)=3*0=0